An insurance salesperson sells an average of 3 life insurance policies per week. Suppose insurance policies sales occur according to a Poisson distribution.

a) What is the probability they will sell 3 or more policies in four weeks?

b) What is the expected value of policies sold in four weeks?

c) What is the variance in the expected number of policies sold?
Answer #1

X ~ Poi (\lambda)

Where \lambda = 3

P(X) = e-\lambda * \lambda X / X!

a)

P(X >= 3) = 1 - P(X <= 2)

= 1 - [ P(X = 0) + P(X = 1) + P(X = 2) ]

= 1 - [ e-3 + e-3 * 3 + e-3 * 32 / 2! ]

= 0.5768

b)

For 4 weeks , \lambda t = 4 * 3 = 12

E(X) = \lambda t = 12

c)

Var(X) =\lambda = 3